Run Plan for week of beginning on Nov 7 2005 updated by Ryan Bayes Nov 10, 2005 This largely resulted from the discussions in the run planning meeting of Nov 7. Note that the times are approximate. /---------------------------------------------------------------- Monday Nov7 evening -finish bulge testing started in the afternoon -check the stopping distribution for the nominal settings and four different settings of B2 (ie B2_nominal, +/-3, +/-6) -purpose of steering the beam into the fringe field to test the direct effects on the beam spot -tests whether the nominal beam is parallel to the spectrometer axis by watching the change in the stopping position as a function of B2 -must be done with constant muon gas degrader setting -find muon degrader setting where stopping distribution is at target (magic number muon_restricted_z = -0.7) while the B2 setting is set at a maximal misalignment (ie B2+6 G) -start with muon degrader at 100% He and bring the restricted muon stopping position back down to -0.7 cm -run at these settings through the night -tests whether we can produce a data set at these settings to test depolarisations due to beam misalignment /---------------------------------------------------------------- Tuesday Nov8 morning -TEC data for quads off, 2.0 T, Be target -Q6 & Q7 steering tests with other quads off -TEC data and rates for graphite target (OPS/JMP) -TEC data and rates for 2004 slit settings, B2 etc. => with graphite if possible evening -29.6+/-1% TEC data @ 2T -purpose of testing the momentum dependence of the TEC efficiency -setup (in Tuesday evening shift?) the 120MeV/c pion tune -t0s with 2.0T field on -4 runs each of PU and MLT triggers -check documentation (on the run page under "Nominal Settings" closely regarding the settings required for the pion tune. Make sure that the beamline settings are correct before letting the pions into the channel /----------------------------------------------------------------------- Wednesday Nov9 graveyard shift -finish the t0 runs if not completed -reproduce and refine the B2 sweep done by Mina on Monday with a few differences. -alter the CO2 percentage in the gas degrader so that the mean restricted stopping position is upstream of the target. This will give us a greater sensitivity to small changes in the stopping position. -use the TEC to determine the upper and lower limits on the maximum and minimum B2 values that can still appear in the TEC -take more data points (nominal, maximum B2, minimum B2, > 4 points to fill out the range at regular intervals). Take the data points at "random" values of B2 during the sweep so changes in atmospheric conditions have a "lesser" effect on the interpretation of the results (Note that the data collected from this is sparse and will need to be repeated on Monday after the TEC is replaced) during the day (maintenance day) Remove and reconstruct the TEC (TEC will not be replaced until Monday) when beam returns -reestablish settings for a data set with the TEC in using, the maximum B2 setting used in the sweep completed on Wednesday overnight shift (for lack of knowledge call this B2 + 6G) -set gas degrader to again center the stopping distribution at -0.7 cm -set slits to configuration established by Makoto at the beginning of the TEC in run /------------------------------------------------------------------------- Thursday Nov10 -collect data for B2 + 6 G data set /------------------------------------------------------------------------- Friday Nov11 Rememberance day Continue collecting B2 + 6 G /------------------------------------------------------------------------ Saturday Nov12 finish collecting B2 + 6G data set -change the B2 setting to the minimum value (found to be -10 G) that still appears in the TEC. -maintain the same slit settings -set gas degrader to maintain the stopping distribution at -0.7 cm start the B2 - 10G data set /------------------------------------------------------------------------ Sunday Nov13 continue the B2 - 10G data set /------------------------------------------------------------------------ Monday Nov14 finish collecting B2 - 6G data set (in the morning) -Collect detailed B2 sweep data Ramp solenoid to 0T -t0 data with 120 MeV/c pion tune, field off -4 runs each of PU and MLT triggers -check documentation closely regarding the settings required for the pion tune. Make sure that the beamline settings are correct before letting the pions into the channel